ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 99, 1997

Abrams, G. Non-induced isomorphisms of matrix rings  343

Blanchard, F. and A. Maass Dynamical properties of expansive one-sided cellular automata

149

Chen, B.-Y. Interaction of Legendre curves and Lagrangian submanifolds

69

Galeeva, R., M. Martens and C. Tresser Inducing, slopes, and conjugacy classes

123

Girardi, M. and W. B. Johnson Universal non-completely-continuous operators

207

Gitik, M. Some results on the nonstationary ideal II  175
Hillman, J. A. On L^{2}-homology and asphericity  271

Johnson, W. B. See M. Girardi

Lawler, G. F. and E. E. Puckette The disconnection exponent for simple random walk

109

Ling, S. On the Q-rational cuspidal subgroup and the component

group of J_{0}(p^{r}) 29

Liu, K. and S. Shelah Cofinalities of elementary substructures of structures on ℵ_{ω}

189

Maass, A. See F. Blanchard

Martens, M. See R. Galeeva

Monterde, J. and O. A. Sánchez-Valenzuela Graded metrics adapted to splittings

231

Navarro-Pascual, J. C. Extreme points and retractions in Banach spaces

335

Parry, W. Cohomology of permutative cellular automata  315

Puckette, E. E. See G. F. Lawler

Qian, N. and R. J. Zimmer Entropy rigidity for semisimple group actions

55

Rigoli, M., M. Salvatori and M. Vignati Subharmonic functions on graphs

1

Rosenmann, A. The normalized cyclomatic quotient associated with presentations of finitely generated groups

285

Salvatori, M. See M. Rigoli Sánchez-Valenzuela, O. A. See J. Monterde

Schindler, R.-D. On a Chang conjecture  221

Shelah, S. See K. Liu Tresser, C. See R. Galeeva Vignati, M. See M. Rigoli Zimmer, R. J. See N. Qian